P/E AND STOCK PRICE Ferrell Inc. recently reported net income of $2 million. It has 260,000 shares of common stock, which currently trades at $23 a share. Ferrell continues to expand and anticipates that 1 year from now, its net income will be $3.1 million. Over the next year, it also anticipates issuing an additional 91,000 shares of stock so that 1 year from now it will have 351,000 shares of common stock. Assuming Ferrells price/earnings ratio remains at its current level, what will be its stock price 1 year from now? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

ANSWER - STOCK PRICE 1 YEAR FROM NOW = $26.41

Market value per share = $23

Net income = $2,000,000

Current no. of shares = 260,000

Earning per share = Net income / Current no. of shares = $2,000,000 / 260,000 = $7.6923076923

Market value per share Earning per share P/E ratio

P / E ratio 23 7.6923076923

P/E ratio 2.99

ONE YEAR FROM NOW

P/E ratio = 2.99

No. of shares = 351,000

Net income = $3,100,000

Earning per share = Net income / Current no. of shares = $3,100,000 / 351,000 = $8.8319088319

Market value per share Earning per share P/E ratio

Market value per share 8,8319088319 2.99 -

Market value per share-26.41